How to I calculate the data?
Plotting Points, Interpreting Data
The South Carolina Statistical Abstract reports the rate of Aggravated Assault (per 10,000 residents) as follows:
YEAR 1985 1986 1987 1988 1989 1990

IV. Assume that the population of South Carolina was 3.1 million in 1980, and grew at a steady 2.5% rate throughout the 80's. (a) Find the S.C. population for the years starting 1985 and ending in 1990, and (b) use those population values to calculate the numbers of aggravated assaults in S.C. during those years. Organize your information in a meaningful format in Excel, and construct a plot of your new data.

Answer by stanbon(75887)

Assume that the population of South Carolina was 3.1 million in 1980, and grew at a steady 2.5% rate throughout the 80's.
(a) Find the S.C. population for the years starting 1985 and ending in 1990,
N(1985) = 3.1(1.025)^5 = 3.507...million
N(1990) = 3.1(1.025)^10 = 3.96826... million
--------------------------------------------------------

(b) use those population values to calculate the numbers of aggravated assaults in S.C. during those years.
Your data for aggravated assaults did not appear in your posting.
==================
Cheers,
Stan H.